OpenStudy (anonymous):

Find the exact value of x if 8^x = "i'll show you below"

OpenStudy (anonymous):

\[8^x = \sqrt{2\sqrt{8\sqrt{2}}}\] Complicated, i know, but if you can solve it WITHOUT using a calculator, pls show working, thanks :P

sam (.sam.):

\[8^{x}=\sqrt{2\sqrt{8(2)^{1/2}}}\] \[8^{x}=\sqrt{2\sqrt{2^{3}(2)^{1/2}}}\] \[8^{x}=\sqrt{2\sqrt{2^{3}(2)^{1/2}}}\] \[8^{x}=\sqrt{2\sqrt{(2)^{7/2}}}\] \[8^{x}=\sqrt{2((2)^{7/2})^{1/2}}\] \[8^{x}=\sqrt{2((2)^{7/4})}\] \[8^{x}=\sqrt{((2)^{11/4})}\] \[2^{3x}=((2)^{11/4})^{1/2}\] \[2^{3x}=((2)^{11/8})\] 3x=11/8 x=11/24

OpenStudy (anonymous):

HOLY MOTHER OF GOD OMFG THATS RIGHT YOU, ARE AWESOME :P sorry, i've been dying over that question for days, and my teacher was refusing to tell me how to do it :P Thanks :)

sam (.sam.):

LOL